Carrington Alarntic Inc. produces motors for yachts. Sales have been very up and down, with some months showing a loss. The company's contribution-format income statement for the most recent month is given below:

                        Sales (20,000 units at $15 per unit)                            $300,000

                        Variable expenses                                                         200,000

                        Contribution margin (CM)                                            100,000

Fixed expenses                                                            120,000

Net operating loss                                                      $ (20,000)

Required:

  1. Compute the company's break-even point in both units and dollars.
  2. The sales manager feels that a $30,000 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in a $150,000 increase in monthly sales. If the sales manager is right, what will be the effect on the company's monthly net operating income or loss?
  3. Refer to the original data. The president is convinced that a 20% reduction in the selling price, combined with an increase of $20,000 in the monthly advertising budget, will cause unit sales to double. What will the new contribution-format income statement look like if these changes are adopted?
  4. Refer to the original data. The company's advertising agency thinks that a new package would help sales. The new package being proposed would increase packaging costs by $1.00 per unit. Assuming no other changes, how many units would have to be sold each month to earn a pretax profit of $20,000?
Answer #1

Break even point in units = Fixed costs/Contribution Margin per unit

= 120,000/5

= 24,000 units

In Dollars = 24000*15 = $360,000

b.Effect on Income = Additional contribution margin- Additional fixed cost

= 150,000*5/15 – 30,000

= $20,000

i.e. income will increase by $20,000

c.

Contribution Margin format Income Statement

Sales

       480,000

Variable Expenses

       400,000

Contribution Margin

          80,000

Fixed expenses

       140,000

Net operating Income

        (60,000)

d.Desired Units = (120,000+20,000)/(15-10-1)

= 35,000 units